OpenStudy (christos):

Please help me factor 8^2-6z-9

hartnn (hartnn):

is it 8z^2 as 1st term ?

OpenStudy (anonymous):

the answer and steps are in the link.

OpenStudy (anonymous):

look for factors of 8 and 9 8=2*4 9=3*3 so: (4z+3)(2z-3)

hartnn (hartnn):

if u wanna try to do it on your own, do you know how to split the middle term for doing factorization ? i assume the question is : \(8z^2-6z-9\)

Parth (parthkohli):

Hello! Yes, these people are correct.

Parth (parthkohli):

The question cannot be \(8^2 - 6z - 9\).

OpenStudy (chmvijay):

yaaa question cant be just 8^2